# Multiple ergodic averages along functions from a Hardy field: convergence, recurrence and combinatorial applications

### v2

@@ 1.3.   Illustrative examples and counterexamples
While part 1 of B holds for any functions $f_{1},\dots,f_{k}\in\mathcal{H}$ of polynomial growth (with an appropriate $W$), parts 2 and 3 require some additional constraints. This is of course unavoidable since for certain functions from $\mathcal{H}$ the conclusions of parts 2 and 3 of B are known to fail. In this subsection we collect examples that illustrate that under some ostensibly natural weakenings/modifications of its conditions, parts 2 and 3 of B are no longer true.
As was discussed in 1.9, if condition 1 of A holds, then the conclusion remains valid when the floor function $\lfloor$.$\rfloor$ . ⌋ is replaced with either the ceiling function $\lceil$.$\rceil$ . ⌉ or the closest integer function [.][.][ . ]. However, as the next example demonstrates, the same is not true in general for condition 2 of A.
@@
Since $f_{1},f_{2}$ satisfy condition 2 of A, it follows that any set $E\subset\mathbb{N}$ of positive upper density contains a triple of the form $\{a,\,a+[f_{1}(n)],a+[f_{2}(n)]\}$. However, the set $E=2\mathbb{N}-1$ does not contain any configurations of the form $\{a,\,a+\lfloor f_{1}(n)\rfloor,a+\lfloor f_{2}(n)\rfloor\}$ or $\{a,\,a+\lceil f_{1}(n)\rceil,a+\lceil f_{2}(n)\rceil\}$.
One may wonder if in A one can relax condition 1 by replacing $\mathrm{span}^{*}(f_{1},\ldots,f_{k})$ with $\mathrm{span}^{*}_{\mathbb{Z}}(f_{1},\ldots,f_{k})=\{c_{1}f_{1}(t)+\ldots+c_{k}f_{k}(t):c_{1},\ldots,c_{k}\in\mathbb{Z}\}$. However, the following example shows that this is not possible.
@@
To see why this is the case, observe that $\{a,\,a+[f_{1}(n)]\}\subset E$ implies that the number $[f_{1}(n)]\alpha$ is within $\frac{1}{8}$ of an integer, which when combined with the condition $|\alpha|\leqslant 1/8$ implies that $f_{1}(n)\alpha$ is within $\frac{1}{4}$ of an integer. Similarly, $\{a,\,a+[f_{2}(n)]\}\subset E$ implies that $f_{2}(n)\beta$ is within $\frac{1}{4}$ of an integer. On the other hand, due to the choice of $f_{1}$ and $f_{2}$, for every $n\in\mathbb{N}$ we have $f_{1}(n)\alpha+f_{2}(n)\beta\equiv 1/2\bmod 1$; so which precludes that $\{a,\,a+[f_{1}(n)],a+[f_{2}(n)]\}\subset E$.
@@
$R=\Big{\{}n\in\mathbb{N}:\mu\big{(}A\cap T^{-[f_{1}(n)]}A\cap\ldots\cap T^{-[f_{k}(n)]}A\big{)}>0\Big{\}}.$ (1.11)
[…]
@@
[…]
In view of 1.17 one might suspect that polynomials with irrational coefficients are the only obstruction to an affirmative answer to 1.16. The following example shows that this is not the case. In fact, no integer linear combination of the functions $f_{1}$ and $f_{2}$ in the following example are in $\mathbb{R}[x]\backslash\mathbb{Z}[x]$.
@@
Let $f_{1}(n)=n+\sqrt{n}$ and $f_{2}(n)=n-\sqrt{n}$. Since the parity of $\lfloor f_{1}(n)\rfloor$ and of $\lfloor f_{2}(n)\rfloor$ are different whenever $n$ is not a perfect square, we see that, for a rotation on two points (i.e. with the same $T$ and $A$ as in 1.17) the set (1.11) has 00 Banach upper density. However, it follows from 1.2 that the set (1.11) has an infinite intersection with the set of perfect squares, and hence is itself infinite.
@@ 1.4.   Outline of the paper
After presenting some preliminaries in Section 2, in Section 3 we briefly explain how each result mentioned in the introduction follows from B. Sections 4 and 5 are dedicated to the proof of B. In Section 4 we show that the nilfactors are characteristic for the expressions involved, and in Section 5 we use this to reduce B to the case of nilsystems, which can then be studied directly using equidistribution results developed in [Ric23]. Finally, in Section 6, we formulate some natural open questions.
Acknowledgements.
We thank N. Frantzikinakis for providing helpful comments on an earlier version of this paper.
@@ 2.1.   Preliminaries on nilsystems and nilmanifolds
2.2.   Preliminaries on Hardy fields
@@
holds for every continuous function $F\in C(X)$. This notion extends the classical notion of uniform distribution mod $\bmod\,1$, which corresponds to the case where $W(N)=N$, the nilmanifold $X=\mathbb{R}/\mathbb{Z}$ is the one dimensional torus and $\nu$ is the Lebesgue measure.
@@
Let $\mathcal{H}$ be a Hardy field and assume $f_{1},\ldots,f_{k}\in\mathcal{H}$ have polynomial growth. Then there exists $W\in\mathcal{H}$ with $1\prec W(t)\ll t$ such that $f_{1},\ldots,f_{k}$ satisfy property 1.
@@
We use induction on $k$. For the base case of this induction, which corresponds to $k=1$, we distinguish between the cases when $\lim_{t\to\infty}|f_{1}(t)-p(t)|<\infty$ for some $p\in\mathbb{R}[t]$, and when $\lim_{t\to\infty}|f_{1}(t)-p(t)|=\infty$ for all $p\in\mathbb{R}[t]$ (noting that the limit must exist since any Hardy field function is eventually monotone.). If we are in the first case then simply take =$\mathcal{I}=\{1\}$, $\mathcal{J}=\emptyset$, $c=\lim_{t\to\infty}f_{1}(t)-p(t)$, and $p_{1}(t)=p(t)+c$. If the latter holds, then we pick $\mathcal{I}=\emptyset$ and $\mathcal{J}=\{1\}$ and we are done.
Now assume the claim has already been proven for $k-1$. This means that for the collection $\{f_{1},\ldots,f_{k-1}\}$ there exist disjoint $\mathcal{I}^{\prime},\mathcal{J}^{\prime}\subset\{1,\ldots,k-1\}$ with $\mathcal{I}^{\prime}\cup\mathcal{J}^{\prime}=\{1,\ldots,k-1\}$, $\{\lambda_{i,j}:i\in\mathcal{I}^{\prime},j\in\mathcal{J}^{\prime}\}\subset\mathbb{R}$, and $\{p_{i}:i\in\mathcal{I}^{\prime}\}\subset\mathrm{poly}(f_{1},\dots,f_{k-1})$ such that conditions 1 and 2 are satisfied. We again distinguish two cases. The first case is when there exists $\{\eta_{j}:j\in\mathcal{J}^{\prime}\}\subset\mathbb{R}$, $\eta_{k}\in\mathbb{R}$, and $p\in\mathbb{R}[t]$ such that $\lim_{t\to\infty}|\sum_{j\in\mathcal{J}^{\prime}}\eta_{j}f_{j}(t)+\eta_{k}f_{k}(t)-p(t)|<\infty$. If we are in this case, then we take =$\mathcal{I}=\mathcal{I}^{\prime}\cup\{k\}$, $\mathcal{J}=\mathcal{J}^{\prime}$, $c=\lim_{t\to\infty}\sum_{j\in\mathcal{J}^{\prime}}\eta_{j}f_{j}(t)+\eta_{k}f_{k}(t)-p(t)$, $p_{k}(t)=\eta_{k}^{-1}(p(t)+c)$, and $\lambda_{k,j}=-\eta_{k}^{-1}\eta_{j}$ for all $j\in\mathcal{J}^{\prime}$. It is then straightforward to check that conditions 1 and 2 are satisfied. The second case is when $\lim_{t\to\infty}|\sum_{j\in\mathcal{J}^{\prime}}\eta_{j}f_{j}(t)+\eta_{k}f_{k}(t)-p(t)|=\infty$ for all $\{\eta_{j}:j\in\mathcal{J}^{\prime}\}\subset\mathbb{R}$, $\eta_{k}\in\mathbb{R}$, and $p\in\mathbb{R}[t]$. But this just means that if we choose $\mathcal{I}=\mathcal{I}^{\prime}$ and $\mathcal{J}=\mathcal{J}^{\prime}\cup\{k\}$ then conditions 1 and 2 hold. ∎
We also need the following corollary of 2.3.
@@
Let $f\in\mathcal{H}$ and assume $f$ satisfies property 1 with respect to some $W\in\mathcal{H}$ with $1\prec W(t)\ll t$. If
@@
Since $f$ satisfies property 1, we either have $|f(t)-p(t)|\ll 1$ for some $p\in\mathbb{R}[t]$ or $|f(t)-p(t)|\succ\log(W(t))$ for all $p\in\mathbb{R}[t]$. If we are in the latter case then we claim that $f(n)$ is uniformly distributed mod $1$ with respect to $W$-averages, which makes (2.2) impossible. To prove the claim, let $\ell$ be the smallest integer for which $f(t)\ll t^{\ell}$. If $\ell>1$ then a standard inductive argument using the version of van der Corput’s trick [^7] Strictly speaking, Corollary 2.6 in [BMR20] is formulated for well distribution (an amplified variant of uniform distribution), but the exact same derivation (using [BMR20, Proposition 2.5]) holds for the version we use here. in [BMR20, Corollary 2.6] reduces the claim to the case $\ell=1$. For $\ell=1$ the claim follows by 2.3.
@@ 3.   Proofs of the corollaries
In this section we explain how all the results in the introduction can be derived from B.
Corollary B1 follows directly from part 1 of B by taking $W(t)=t$ and Corollary B2 follows directly from part 2 of B in the case of an ergodic system. Corollaries B3 and B4 follow from parts 2 and 3, respectively, together with the fact that for any bounded sequence $a:\mathbb{N}\to\mathbb{R}$ taking non-negative values we have888 For completeness, we provide a proof of (3.1): Let $c_{N}(M)=\frac{M}{W(N)}\left(w(M)-w(M+1)\right)$ if $M<N$, and $c_{N}(N)=\frac{Nw(N)}{W(N)}$. Since $w(n)$ is eventually non-increasing, $c_{N}(M)\geqslant 0$ for sufficiently large $M$. The following identity, which can be readily checked by induction, shows that the weighted averages of $a$ are related to the unweighted averages via the coefficients $c_{N}(M)$: $\frac{1}{W(N)}\sum_{n=1}^{N}w(n)a(n)=\sum_{M=1}^{N}c_{N}(M)\biggl{(}\frac{1}{M}\sum_{n=1}^{M}a(n)\biggr{)}.$ Applying this formula with $a\equiv 1$ it follows that, for each $N\in\mathbb{N}$, the coefficients $c_{N}(1),\dots,c_{N}(N)$ add up to $1$. Therefore, for an arbitrary non-negative $a:\mathbb{N}\to\mathbb{R}$, for any $N$ there exists some $M=M(N)\leqslant N$ such that $\frac{1}{M}\sum_{n=1}^{M}a(n)\geqslant\frac{1}{W(N)}\sum_{n=1}^{N}w(n)a(n).$ Moreover, since $c_{N}(M)\to 0$ as $N\to\infty$ for any fixed $N$, we have that $M(N)\to\infty$ as $N\to\infty$, so we get (3.1).
@@
A is obtained by combining Corollaries B3 and B4 with Furstenberg’s correspondence principle (1.11).
@@
Since all the $f_{i}$ have polynomial growth, there is a maximum degree $d$ among all polynomials in $\mathrm{poly}(f_{1},\dots,f_{k})$. If all $p\in\mathrm{poly}(f_{1},\dots,f_{k})$ satisfies $p(0)=0$, then $\mathrm{poly}(f_{1},\dots,f_{k})\subset\mathrm{span}(t,t^{2},\dots,t^{d})$. Since the polynomials $t,t^{2},\dots,t^{d}$ are jointly intersective, the result follows from part 2 of A. ∎
@@
[…]
@@
Corollary A3 can be shown to follow from Corollary A1. Alternative, since ${\mathcal{K}}={\mathcal{K}}({\mathcal{L}},q)$ where $q(t)=t$ and =${\mathcal{L}}=\{t^{c}\log_{m}^{r}:c\in(0,1),m\in\mathbb{N},r>0\}$, by using 1.10 one can derive Corollary A3 from part 2 of A.
The following theorem provides a convenient description of functions which satisfy Condition 2 in the A and, in particular, implies 1.10.
@@
Then functions $f_{1},\dots,f_{k}$ from $\mathcal{H}$ satisfy Condition 2 of A if and only if $\{f_{1},\dots,f_{k}\}\subset{\mathcal{K}}({\mathcal{L}},q)$ for some ${\mathcal{L}}$ and $q$ as above.
@@
Suppose that $f_{1},\dots,f_{k}\in{\mathcal{K}}({\mathcal{L}},q)$ for some ${\mathcal{L}}$ and $q$ as in the statement of 3.1. Then $\mathrm{poly}(f_{1},\dots,f_{k})\subset q\mathbb{R}[t]$, which is spanned by the jointly intersective collection $\{t^{\ell}q(t):\ell\geqslant 0\}$. This shows that $f_{1},\dots,f_{k}\in\mathcal{H}$ satisfy Condition 2 of A.
Conversely, let $\mathcal{H}$ be a Hardy field and let $f_{1},\dots,f_{k}\in\mathcal{H}$ satisfy Condition 2 of A. Using [Ric23, Lemma A.3] we can find $g_{1},\dots,g_{m}\in\mathrm{span}^{*}(f_{1},\dots,f_{k})$ of different growth satisfying $t^{\ell-1}\prec g_{j}(t)\prec t^{\ell}$ for all $j=1,\dots,m$ and some $\ell\in\mathbb{N}$ (which depends on $j$) and $p_{1},\dots,p_{s}\in\mathrm{poly}(f_{1},\dots,f_{k})$ such that for every $i=1,\dots,k$ there exists $f_{i}^{*}\in\mathrm{span}^{*}(g_{1},\dots,g_{m},p_{1},\dots,p_{s})$ with $\lim|f_{i}(t)-f_{i}^{*}(t)|=0$. Condition 2 implies that $p_{1},\dots,p_{s}\in\mathrm{span}(q_{1},\dots,q_{\ell})$ for some jointly intersective polynomials $q_{1},\dots,q_{\ell}\in\mathbb{Z}[t]$. In view of [BLL08, Proposition 6.1] there exists an intersective polynomial $q$ such that $q_{1},\dots,q_{\ell}\in q\mathbb{Z}[t]$, and hence $p_{1},\dots,p_{s}\in q\mathbb{R}[t]$. Letting =${\mathcal{L}}=\{g_{1},\dots,g_{m}\}$, we conclude that $f_{1},\dots,f_{k}\in{\mathcal{K}}({\mathcal{L}},q)$.
@@
In view of A, it suffices to show that $F_{\ell}$ satisfies condition 1 in that theorem. Considering a finite Taylor expansion of $f_{j}$, it follows that each function in $F_{\ell}$ “almost” belongs to -$\nabla\text{-}\mathrm{span}(f_{1},\ldots,f_{k})$, in the sense that for any $f\in F_{\ell}$ there exists $g\in\nabla\text{-}\mathrm{span}(f_{1},\ldots,f_{k})$ such that $|f(n)-g(n)|\to 0$. More generally, for any $f\in\mathrm{span}^{*}(F_{\ell})$ there exists $g\in\nabla\text{-}\mathrm{span}(f_{1},\ldots,f_{k})$ such that $|f(n)-g(n)|\to 0$.

### v3

@@ 1.3.  Illustrative examples and counterexamples
While part (i) of Theorem B holds for any functions $f_{1},\dots,f_{k}\in\mathcal{H}$ of polynomial growth (with an appropriate $W$), parts (ii) and (iii) require some additional constraints. This is of course unavoidable since for certain functions from $\mathcal{H}$ the conclusions of parts (ii) and (iii) of Theorem B are known to fail. In this subsection we collect examples that illustrate that under some ostensibly natural weakenings/modifications of its conditions, parts (ii) and (iii) of Theorem B are no longer true.
As was discussed in ˜1.9, if condition (1) of Theorem A holds, then the conclusion remains valid when the floor function $\lfloor.\rfloor$ is replaced with either the ceiling function $\lceil.\rceil$ or the closest integer function $[.]$. However, as the next example demonstrates, the same is not true in general for condition (2) of Theorem A.
@@
Since $f_{1},f_{2}$ satisfy condition (2) of Theorem A, it follows that any set $E\subset\mathbb{N}$ of positive upper density contains a triple of the form $\{a,\,a+[f_{1}(n)],a+[f_{2}(n)]\}$. However, the set $E=2\mathbb{N}-1$ does not contain any configurations of the form $\{a,\,a+\lfloor f_{1}(n)\rfloor,a+\lfloor f_{2}(n)\rfloor\}$ or $\{a,\,a+\lceil f_{1}(n)\rceil,a+\lceil f_{2}(n)\rceil\}$.
One may wonder if in Theorem A one can relax condition (1) by replacing $\mathrm{span}^{*}(f_{1},\ldots,f_{k})$ with $\mathrm{span}^{*}_{\mathbb{Z}}(f_{1},\ldots,f_{k})=\{c_{1}f_{1}(t)+\ldots+c_{k}f_{k}(t):c_{1},\ldots,c_{k}\in\mathbb{Z}\}$. However, the following example shows that this is not possible.
@@
To see why this is the case, observe that $\{a,\,a+[f_{1}(n)]\}\subset E$ implies that the number $f_{1}(n)]\alpha$ is within $\frac{1}{8}$ of an integer, which when combined with the condition $|\alpha|\leqslant 1/8$ implies that $f_{1}(n)\alpha$ is within $\frac{1}{4}$ of an integer. Similarly, $\{a,\,a+[f_{2}(n)]\}\subset E$ implies that $f_{2}(n)\beta$ is within $\frac{1}{4}$ of an integer. On the other hand, due to the choice of $f_{1}$ and $f_{2}$, for every $n\in\mathbb{N}$ we have $f_{1}(n)\alpha+f_{2}(n)\beta\equiv 1/2\bmod 1$; so which precludes that $\{a,\,a+[f_{1}(n)],a+[f_{2}(n)]\}\subset E$.
@@
$R=\Big\{n\in\mathbb{N}:\mu\big(A\cap T^{-[f_{1}(n)]}A\cap\ldots\cap T^{-[f_{k}(n)]}A\big)>0\Big\},$ (1.11)
is either finite or satisfies $\bar{d}(R)>0$ (or at least $d^{*}(R)>0$ [^4] Given a set $R\subset\mathbb{N}$ the upper Banach density of $R$ is defined by $d^{*}(R):=\lim_{N\to\infty}\sup_{M\in\mathbb{N}}\frac{1}{N}|R\cap\{M,M+1,\dots,M+N\}|.$ )?
@@
Examples analogous to ˜1.17 exist where $\cdot$] is replaced by $\lfloor\cdot\rfloor$ and $\lceil\cdot\rceil$. One can also create an example where the dynamical system is an irrational rotation instead of a rotation on only two points. Finally we point out that, by choosing $\alpha$ appropriately one can make the set $\big\{n\in\mathbb{N}:\{n\alpha-1/2\}<C/n\big\}$ be arbitrarily sparse.
In view of ˜1.17 one might suspect that polynomials with irrational coefficients are the only obstruction to an affirmative answer to ˜1.16. The following example shows that this is not the case. In fact, no integer linear combination of the functions $f_{1}$ and $f_{2}$ in the following example belongs to $\mathbb{R}[x]\backslash\mathbb{Z}[x]$.
@@
Let $f_{1}(n)=n+\sqrt{n}$ and $f_{2}(n)=n-\sqrt{n}$. Since $\lfloor f_{1}(n)\rfloor$ and $\lfloor f_{2}(n)\rfloor$ have different parity whenever $n$ is not a perfect square, we see that, for a rotation on two points (i.e. with the same $T$ and $A$ as in ˜1.17) the set (1.11) has 0 Banach upper density. However, it follows from ˜1.2 that the set (1.11) has an infinite intersection with the set of perfect squares, and hence is itself infinite.
@@ 1.4.  Outline of the paper
After presenting some preliminaries in Section˜2, in Section˜3 we briefly explain how each result mentioned in the introduction follows from Theorem B. Sections 4 and 5 are dedicated to the proof of Theorem B. In Section˜4 we show that the nilfactors are characteristic for the expressions involved, and in Section˜5 we use this to reduce Theorem B to the case of nilsystems, which can then be studied directly using equidistribution results developed in [Ric23]. Finally, in Section˜6, we formulate some natural open questions.
Acknowledgments.
We thank N. Frantzikinakis for providing helpful comments on an earlier version of this paper. We thank Saúl Rodríguez Martín for pointing out an error in the definition of -$\nabla\text{-}\mathrm{span}(f_{1},\ldots,f_{k})$ in the version of this paper published in Advances in Mathematics; the error has been corrected in the present arXiv version.
@@ 2.1.  Preliminaries on nilsystems and nilmanifolds
$2.$  Preliminaries on Hardy fields
@@
holds for every continuous function $F\in C(X)$. This notion extends the classical notion of uniform distribution $\bmod\,1$, which corresponds to the case where $W(N)=N$, the nilmanifold $X=\mathbb{R}/\mathbb{Z}$ is the one dimensional torus and $\nu$ is the Lebesgue measure.
@@
Let $\mathcal{H}$ be a Hardy field and assume $f_{1},\ldots,f_{k}\in\mathcal{H}$ have polynomial growth. Then there exists $W\in\mathcal{H}$ with $1\prec W(t)\ll t$ such that $f_{1},\ldots,f_{k}$ satisfy property (P).
@@
We use induction on $k$. For the base case of this induction, which corresponds to $k=1$, we distinguish between the cases when $\lim_{t\to\infty}|f_{1}(t)-p(t)|<\infty$ for some $p\in\mathbb{R}[t]$, and when $\lim_{t\to\infty}|f_{1}(t)-p(t)|=\infty$ for all $p\in\mathbb{R}[t]$ (noting that the limit must exist since any Hardy field function is eventually monotone.). If we are in the first case then simply take $\mathcal{I}=\{1\}$, $\mathcal{J}=\emptyset$, $c=\lim_{t\to\infty}f_{1}(t)-p(t)$, and $p_{1}(t)=p(t)+c$. If the latter holds, then we pick $\mathcal{I}=\emptyset$ and $\mathcal{J}=\{1\}$ and we are done.
Now assume the claim has already been proven for k1k-1. This means that for the collection $\{f_{1},\ldots,f_{k-1}\}$ there exist disjoint $\mathcal{I}^{\prime},\mathcal{J}^{\prime}\subset\{1,\ldots,k-1\}$ with $\mathcal{I}^{\prime}\cup\mathcal{J}^{\prime}=\{1,\ldots,k-1\}$, $\{\lambda_{i,j}:i\in\mathcal{I}^{\prime},j\in\mathcal{J}^{\prime}\}\subset\mathbb{R}$, and $\{p_{i}:i\in\mathcal{I}^{\prime}\}\subset\mathrm{poly}(f_{1},\dots,f_{k-1})$ such that conditions (a) and (b) are satisfied. We again distinguish two cases. The first case is when there exists $\{\eta_{j}:j\in\mathcal{J}^{\prime}\}\subset\mathbb{R}$, $\eta_{k}\in\mathbb{R}$, and $p\in\mathbb{R}[t]$ such that $\lim_{t\to\infty}|\sum_{j\in\mathcal{J}^{\prime}}\eta_{j}f_{j}(t)+\eta_{k}f_{k}(t)-p(t)|<\infty$. If we are in this case, then we take $\mathcal{I}=\mathcal{I}^{\prime}\cup\{k\}$, $\mathcal{J}=\mathcal{J}^{\prime}$, $c=\lim_{t\to\infty}\sum_{j\in\mathcal{J}^{\prime}}\eta_{j}f_{j}(t)+\eta_{k}f_{k}(t)-p(t)$, $p_{k}(t)=\eta_{k}^{-1}(p(t)+c)$, and $\lambda_{k,j}=-\eta_{k}^{-1}\eta_{j}$ for all $j\in\mathcal{J}^{\prime}$. It is then straightforward to check that conditions (a) and (b) are satisfied. The second case is when $\lim_{t\to\infty}|\sum_{j\in\mathcal{J}^{\prime}}\eta_{j}f_{j}(t)+\eta_{k}f_{k}(t)-p(t)|=\infty$ for all $\{\eta_{j}:j\in\mathcal{J}^{\prime}\}\subset\mathbb{R}$, $\eta_{k}\in\mathbb{R}$, and $p\in\mathbb{R}[t]$. But this just means that if we choose $\mathcal{I}=\mathcal{I}^{\prime}$ and $\mathcal{J}=\mathcal{J}^{\prime}\cup\{k\}$ then conditions (a) and (b) hold. ∎
We also need the following corollary of ˜2.3.
@@
Let $f\in\mathcal{H}$ and assume $f$ satisfies property (P) with respect to some $W\in\mathcal{H}$ with $1\prec W(t)\ll t$. If
@@
Since $f$ satisfies property (P), we either have $|f(t)-p(t)|\ll 1$ for some $p\in\mathbb{R}[t]$ or $|f(t)-p(t)|\succ\log(W(t))$ for all $p\in\mathbb{R}[t]$. If we are in the latter case then we claim that $f(n)$ is uniformly distributed mod $1$ with respect to $W$-averages, which makes (2.2) impossible. To prove the claim, let $\ell$ be the smallest integer for which $f(t)\ll t^{\ell}$. If $\ell>1$ then a standard inductive argument using the version of van der Corput’s trick [^7] Strictly speaking, Corollary 2.6 in [BMR20] is formulated for well distribution (an amplified variant of uniform distribution), but the exact same derivation (using [BMR20, Proposition 2.5]) holds for the version we use here. in [BMR20, Corollary 2.6] reduces the claim to the case $\ell=1$. For $\ell=1$ the claim follows by ˜2.3.
@@ 3.  Proofs of the corollaries
In this section we explain how all the results in the introduction can be derived from Theorem B.
Corollary B1 follows directly from part (i) of Theorem B by taking $W(t)=t$ and Corollary B2 follows directly from part (ii) of Theorem B in the case of an ergodic system. Corollaries B3 and B4 follow from parts (ii) and (iii), respectively, together with the fact that for any bounded sequence $a:\mathbb{N}\to\mathbb{R}$ taking non-negative values we have [^8] For completeness, we provide a proof of (3.1): Let $c_{N}(M)=\frac{M}{W(N)}\left(w(M)-w(M+1)\right)$ if $M<N$, and $c_{N}(N)=\frac{Nw(N)}{W(N)}$. Since $w(n)$ is eventually non-increasing, $c_{N}(M)\geqslant 0$ for sufficiently large $M$. The following identity, which can be readily checked by induction, shows that the weighted averages of $a$ are related to the unweighted averages via the coefficients $c_{N}(M)$: $\frac{1}{W(N)}\sum_{n=1}^{N}w(n)a(n)=\sum_{M=1}^{N}c_{N}(M)\biggl(\frac{1}{M}\sum_{n=1}^{M}a(n)\biggr).$ Applying this formula with $a\equiv 1$ it follows that, for each $N\in\mathbb{N}$, the coefficients $c_{N}(1),\dots,c_{N}(N)$ add up to $1$. Therefore, for an arbitrary non-negative $a:\mathbb{N}\to\mathbb{R}$, for any $N$ there exists some $M=M(N)\leqslant N$ such that $\frac{1}{M}\sum_{n=1}^{M}a(n)\geqslant\frac{1}{W(N)}\sum_{n=1}^{N}w(n)a(n).$ Moreover, since $c_{N}(M)\to 0$ as $N\to\infty$ for any fixed $N$, we have that $M(N)\to\infty$ as $N\to\infty$, so we get (3.1).
@@
Theorem A is obtained by combining Corollaries B3 and B4 with Furstenberg’s correspondence principle (˜1.11).
@@
Since all the $f_{i}$ have polynomial growth, there is a maximum degree $d$ among all polynomials in $\mathrm{poly}(f_{1},\dots,f_{k})$. If all $p\in\mathrm{poly}(f_{1},\dots,f_{k})$ satisfies $p(0)=0$, then $\mathrm{poly}(f_{1},\dots,f_{k})\subset\mathrm{span}(t,t^{2},\dots,t^{d})$. Since the polynomials $t,t^{2},\dots,t^{d}$ are jointly intersective, the result follows from part (2) of Theorem A. ∎
@@
The conclusion follows immediately from Corollary A1 when the rounding function is the closest integer function $\cdot$]. We next prove the result for the floor function $\lfloor\cdot\rfloor$, noting that the case for the ceiling function $\lceil\cdot\rceil$ is analogous. Since $\lfloor x\rfloor=[x-1/2]$ and when $n,c\in\mathbb{N}$ we have $\lfloor n^{c}\rfloor=n^{c}=[n^{c}]$, it follows that, for each $1,\dots,k$, $\lfloor n^{c_{i}}\rfloor=[f_{i}(n)]$ where
@@
Corollary A3 can be shown to follow from Corollary A1. Alternative, since ${\mathcal{K}}={\mathcal{K}}({\mathcal{L}},q)$ where $q(t)=t$ and ${\mathcal{L}}=\{t^{c}\log_{m}^{r}:c\in(0,1),m\in\mathbb{N},r>0\}$, by using ˜1.10 one can derive Corollary A3 from part (2) of Theorem A.
The following theorem provides a convenient description of functions which satisfy Condition (2) in the Theorem A and, in particular, implies ˜1.10.
@@
Then functions $f_{1},\dots,f_{k}$ from $\mathcal{H}$ satisfy Condition (2) of Theorem A if and only if $\{f_{1},\dots,f_{k}\}\subset{\mathcal{K}}({\mathcal{L}},q)$ for some $\mathcal{L}$} and $q$ as above.
@@
Suppose that $f_{1},\dots,f_{k}\in{\mathcal{K}}({\mathcal{L}},q)$ for some $\mathcal{L}$} and $q$ as in the statement of ˜3.1. Then $\mathrm{poly}(f_{1},\dots,f_{k})\subset q\mathbb{R}[t]$, which is spanned by the jointly intersective collection $\{t^{\ell}q(t):\ell\geqslant 0\}$. This shows that $f_{1},\dots,f_{k}\in\mathcal{H}$ satisfy Condition (2) of Theorem A.
Conversely, let $\mathcal{H}$ be a Hardy field and let $f_{1},\dots,f_{k}\in\mathcal{H}$ satisfy Condition (2) of Theorem A. Using [Ric23, Lemma A.3] we can find $g_{1},\dots,g_{m}\in\mathrm{span}^{*}(f_{1},\dots,f_{k})$ of different growth satisfying $t^{\ell-1}\prec g_{j}(t)\prec t^{\ell}$ for all $j=1,\dots,m$ and some $\ell\in\mathbb{N}$ (which depends on $j$) and $p_{1},\dots,p_{s}\in\mathrm{poly}(f_{1},\dots,f_{k})$ such that for every $i=1,\dots,k$ there exists $f_{i}^{*}\in\mathrm{span}^{*}(g_{1},\dots,g_{m},p_{1},\dots,p_{s})$ with $\lim|f_{i}(t)-f_{i}^{*}(t)|=0$. Condition (2) implies that $p_{1},\dots,p_{s}\in\mathrm{span}(q_{1},\dots,q_{\ell})$ for some jointly intersective polynomials $q_{1},\dots,q_{\ell}\in\mathbb{Z}[t]$. In view of [BLL08, Proposition 6.1] there exists an intersective polynomial $q$ such that $q_{1},\dots,q_{\ell}\in q\mathbb{Z}[t]$, and hence $p_{1},\dots,p_{s}\in q\mathbb{R}[t]$. Letting ${\mathcal{L}}=\{g_{1},\dots,g_{m}\}$, we conclude that $f_{1},\dots,f_{k}\in{\mathcal{K}}({\mathcal{L}},q)$.
@@
In view of Theorem A, it suffices to show that $F_{\ell}$ satisfies condition (1) in that theorem. Considering a finite Taylor expansion of $f_{j}$, it follows that each function in $F_{\ell}$ “almost” belongs to -$\nabla\text{-}\mathrm{span}(f_{1},\ldots,f_{k})$, in the sense that for any $f\in F_{\ell}$ there exists $g\in\nabla\text{-}\mathrm{span}(f_{1},\ldots,f_{k})$ such that $|f(n)-g(n)|\to 0$. More generally, for any $f\in\mathrm{span}^{*}(F_{\ell})$ there exists $g\in\nabla\text{-}\mathrm{span}(f_{1},\ldots,f_{k})$ such that $|f(n)-g(n)|\to 0$.